OpenStudy (thomas5267):

Invert \(2e^x-\frac{1}{e^x}\).

OpenStudy (thomas5267):

Inverse of \(2y-\frac{1}{y}\).

OpenStudy (dan815):

okay suppose X=2y-1/y

OpenStudy (dan815):

oh u make e^x=y

OpenStudy (dan815):

and no need for that stuff

OpenStudy (dan815):

just say y=2*e^x - 1/e^x solve for x in terms of y

OpenStudy (thomas5267):

I know that but what are the steps? Wolfram Alpha gives \(\log\left(\frac{1}{4}(x-\sqrt{x^2+8}\right)\) but I have no idea where that came from.

OpenStudy (dan815):

okay

OpenStudy (dan815):

lets work it out

OpenStudy (perl):

actually im more comfortable with the substitution u = e^x

OpenStudy (dan815):

hmm

OpenStudy (dan815):